HUBBLE SPACE TELESCOPE DAILY REPORT #927 PERIOD COVERED: 0800 EST Daily Status Report as of 194/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Set of FOC/96 4804 (UV Imaging of Nearby Galaxies with FOC/96) The Faint Object Camera was used to successfully complete two sets of snapshot ultraviolet observations of four nearby galaxies: UGC-7831, UGC-6870, UGC-7898, and UGC-7772. All of the exposures lasted for ten minutes and used filter F220W. The first four observations were executed without apparent problems. A communication stoppage occurred during the last two observations of galaxies ESO-0514-3 and UGC-9179. (See Note 2.4) The status of these observations is unknown at this time. 1.2 Uplinked Updated Ephemeris Tables , Per MOSES 2871 This event was not executed due to the current Safemode event. (See Note 2.4) 1.3 Completed FOS/BL 4952 (QSO Absorption System Snapshot Survey - Part 2 of 3: Cycle 3 High) This observations was not executed due to the current Safemode event. (See Note 2.4) 1.4 Completed FOS/BL 4661 (Ultraviolet Eclipsing Mapping of the Accretion Flow in the Eclipsing Int.) This observations was not executed due to the current Safemode event. (See Note 2.4) 1.5 Completed Two Set of WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) This observations was not executed due to the current Safemode event. (See Note 2.4) 1.6 Completed Two Set of WFC 4381 (HST Medium - Deep Survey - Cycle 3) This observations was not executed due to the current Safemode event. (See Note 2.4) 1.7 Completed WFC PC 4109 (WF/PC Cycle 2 Calibration: Rapid Internal Monitor) This observations was not executed due to the current Safemode event. (See Note 2.4) 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 0 Successful : 0 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 3 Successful : 3 2.3 Significant HSTARs: None 2.4 Operations Note: An on-going SPC load was inadvertently overwritten with new SPC load at 193/1739Z (HSTAR-4218) causing the SPC load processing to idle in a hard wait state. All communications were subsequently lost. Communications were restored at approximately 193/2010Z with reconfiguration to LGAs. Inertial Hold Safemode was commanded at 193/2034Z using COP 6.1. Recovery procedures are in process, with the start of the Health and Safety load scheduled for 194/1600Z, and the return to SMS processing scheduled for 195/2200Z. At the time when the SPC load was overwritten, the FOC was in between observations on Side 2 (f/96) with its high voltage turned on, and the FOS was in low voltage mode on Side 2. The other three instruments were not active. The FOC was safed automatically when Inertial Hold Safemode was commanded. Safing RTCS's are being uplinked as part of the Health and Safety load, and science commanding will resume at 195/2200Z when the SMS begins. DF-224 Acquisition Logic , EPS battery ROC, and battery benchmark SOC patches were successfully installed on day 193. Concomitant EPS data base changes were also implemented. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete PC 5014 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 2) This observations will not executed due to the current Safemode event. (See Note 2.4) 3.2 Complete WFC 4381 (Medium Deep Sky survey - Cycle 3) This observations will not executed due to the current Safemode event. (See Note 2.4) 3.3 Complete FOS/BL 4953 (QSO Absorption System Snapshot Survey - Part 3 of 3 : Cycle 3 High) This observations will not executed due to the current Safemode event. (See Note 2.4) 3.4 Complete Two Sets of FGS 5055 (A Critical Test of the Galactic Escape Velocity at R(Sun): Cycle 1 Obs.) The first set of observations will not executed due to the current Safemode event. (See Note 2.4) **** Return to SMS is scheduled for 195/2200Z. **** 3.5 Complete Two Sets of WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) The first set of observations will not executed due to the current Safemode event. (See Note 2.4) **** Return to SMS is scheduled for 195/2200Z. **** 3.6 Complete FOC/96 4969 (Search for QSOs Suitable for Subsequent Observations of He II ) 3.7 Complete GHRS/2 4476 (Abundance and Distribution of Atomic Deuterium in the Jovian Upper Atmosphere) 3.8 Complete PC 4929 (WF/PC Cycle 2: Improved Dark Calibration) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: The start of the Health and Safety load scheduled for 194/1600Z, and the return to SMS processing scheduled for 195/2200Z. -END-